Presentation is loading. Please wait.

Presentation is loading. Please wait.

December 8 th, 2006Detector Construction status report Yves Déclais 1 OPERA Detector Preparation 1. Electronic Detectors: status of installation 2. Brick.

Similar presentations


Presentation on theme: "December 8 th, 2006Detector Construction status report Yves Déclais 1 OPERA Detector Preparation 1. Electronic Detectors: status of installation 2. Brick."— Presentation transcript:

1 December 8 th, 2006Detector Construction status report Yves Déclais 1 OPERA Detector Preparation 1. Electronic Detectors: status of installation 2. Brick preparation lead production BAM BMS Conclusions about the filling 3. CNGS beam status 4. Perspectives for the run in Conclusions

2 December 8 th, 2006Detector Construction status report Yves Déclais 2 1. Electronic Detectors

3 December 8 th, 2006Detector Construction status report Yves Déclais 3 Electronic detectors installation SM1 Veto BMS Target tracker SM2 Spectrometer: XPC, HPT, RPC, magnet All installed and running except: VETO : commissioning start soon HPT/SM2 : first half of 2007

4 December 8 th, 2006Detector Construction status report Yves Déclais 4 DAQ Event builder Root Alignment Data analysis Brick finding Gateway DB CNGS DB (Oracle) BMM TTdaqRPCdaq HPTdaq Majority trigger Same DAQ for RPC-XPC-VETO 54 sensors X-Y coincidence L-R coincidence N planes coinc? 992 sensors Trigger cross-check? Tests procedure 105 sensors … TTsensors … RPCsensors… HPTsensors ASCII Time coincidence between subdetectors (merge sort) Event classification wrt the beam UTC Event header production Raw data DB postGres (DAQ) Oracle Event DB (Oracle) DAQ manager Beam syncronizer Decodes CNGS early warning Flag the DAQ cycle wrt the beam CNGS Early Warning Connects the DAQs Send flagged cycles Writes GPS dates ASCII ASCII-2-Root

5 December 8 th, 2006Detector Construction status report Yves Déclais 5 2. Brick Preparation - Lead - BAM - BMS

6 6 Westfälische Wilhelms- Universität, Institut für Kernphysik, Münster D. Frekers V. Pilipenko Lead production Oct.-Dec. 2006

7 December 8 th, 2006Detector Construction status report Yves Déclais 7  -activity measured October Edge control 3.Thickness control 4.Status of lead production in October 2006 General situation: We are getting into a much more stable production. Quality control is now routine also among the general personnel at JLGoslar Development in QC

8 December 8 th, 2006Detector Construction status report Yves Déclais 8 sample Date of measurements total5-5.5 MeV total5-5.5 MeVtotal5-5.5 MeV PbCa0.05 (2985), washed 0.75% NGL 5  PbCa0.05 (2987), washed 0.75% NGL 4.5   1. 2 PbCa0.04 (2991), washed 0.75% NGL 6.3   1 PbCa0.04 (2993), washed 0.75% NGL 4.3   0. 3 Results of PbCa0.04 and PbCa0.05  -activity measurements Oct-06

9 December 8 th, 2006Detector Construction status report Yves Déclais 9 More on radioactivity  activity measurements  210 Po : rebuilding from 210 Bi, no migration visible

10 December 8 th, 2006Detector Construction status report Yves Déclais 10 Edge control Control improved by introducing one additional microscope with 4.5 x ocular in OPERA hall at JL Goslar site (see photo).

11 December 8 th, 2006Detector Construction status report Yves Déclais 11

12 December 8 th, 2006Detector Construction status report Yves Déclais 12 Thickness control: We introduced two additional sensors for the thickness control (see photo).

13 December 8 th, 2006Detector Construction status report Yves Déclais 13 Both sensors perform line measurements. If the thickness of the lead is outside specs, an audio alarm signal will be issued and a light will flash. The rolling will stop. Thickness control: flash light

14 14 PbCa0.05 rolling Ca content uOct-2: decision was taken by Collab. Board to increase Ca content to PbCa0.07  We produced 3 charges of the PbCa0.05. This increase of the amount of Ca in the lead gave again difficulties with rolling, i.e. keeping specs was not possible. The decision was then taken by JL Goslar to go back to the PbCa0.04.

15 December 8 th, 2006Detector Construction status report Yves Déclais 15 PbCa0.04 has enough amount of Ca to avoid corrosion on surface, which was also barely observed in case of PbCa0.03. Surface corrosion of the PbCa0.04 is under permanent QC. truck-6, which was stopped due to problem of stickiness and corrosion on plates, has now been scrutinized and re-checked and delivered in November (feed-back so far OK). production has now been reduced to < plates a day until BAM/BMS can ramp up to full performance. next truck delivery foreseen in Jan-07 to avoid production stop, a repository of pallets has been ordered -- ~7000 add‘l pallets will give a total of ~11000 pallets in a buffer (equiv. to plates or 6 trucks) plates per day has been achieved Status of lead production

16 December 8 th, 2006Detector Construction status report Yves Déclais 16 Brick components Lead pallet production BAM fine tuning BAM additional machines Drum filling station BAM Project Status Report S. Buontempo (on behalf of BAM working group) OPERA Emulsion Workshop, Nagoya, December 2006

17 17 Brick components – Nov 06 Al Spider : produced, delivered in Tecnocut Plastic protection : produced and delivered in Tecnocut Assembled spider+plastic protection : 5000 assembled and 3000 delivered in GS Plastic cover : produced, delivered in GS Al tape: all delivered in GS, Skate tape: all delivered in GS CS box: new prototype (20 pcs) produced according to Nagoya requests 5000 being produced now Bar code labels: 2000 delivered in GS, all the rest ordered now. Skate thickness labes: being delivered this week all the rest to be ordered CS box glue: for bricks being delivered this week, all the rest to be ordered

18 18 Brick components – Nov 06 Spider-Plastic protection thermal assembly tool Minor problems of mechanical stability of spider position, turned out to be VERY negative for brick precision. This solution was abandoned and pieces are now being glued. Net result is a longer assembly time (about a factor 2 !)

19 19 Lead Transportation pallets pieces ordered (1000 spare). Delivered (according to Tecnocut). In order to increase the buffer in JL Goslar, Uni Muensterhas ordered additional 6600 now. Delivery in 30 days.

20 20 BAM fine tuning Lead Dispatching line : DONE Lead transportation line: DONE 5 piling pressing stations: DONE (need to refurbish the emulsion storage containers) Al wrapping machine n.1 : mechanically DONE, being registered today Al Wrapping station n.2 : DONE Skate station : DONE CS box station : DONE Data matrix station: DONE brick dimension control : DONE

21 21 BAM additional machines Skate Additional Label : Being delivered in GS in next week (for bricks) Skate additional label machine: being ordered in GS Glue for CS box: Being delivered in GS in next week (for bricks) Glue dispenser for CS box: Being ordered in GS in next week

22 22 Drum station Supposed to be delivered this week in GS: Postponed for mechanical defects:TecnoCut found some problems in the precision of the pusher devices which demand some more fine tuning. This operation is much cheaper to be done in their site than in underground lab. For this reason the shipment has been postponed. Most probably to january 8th as there are difficulties to get the material and the re- machining of the pieces before X-mas.

23 23 BAM production schedule Today and Monday: re-printing of DM on old bricks (with DM at lower level) 250 bricks (mechanical) are already produced (piling-pressing) with light on. Next week they will be completed with Al wrapping and skate + CS with new glue + additional label (if delivered ?) The mechanical quality of these 250 bricks is MUCH better that the first 1000 produced in october. The piling is done with 5 stations with a speed of 50 bricks/hour (10 brick/station/hour). As soon as the 250 bricks are completed we plan to start the real brick production (presumably in second half on next week if all material – glue/label is delivered).  the goal is to reach 3 drums / day in january as a routine speed ( 702 bricks  ~ lead plates)

24 December 8 th, 2006Detector Construction status report Yves Déclais 24 BMS report u Brick filling and extraction n Situation after the October run u BMS commissioning progress n Improvements, solved problems n Mechanical interventions n Filling rate issue n Fence u Brick production restart n Final tuning to be done with final bricks

25 December 8 th, 2006Detector Construction status report Yves Déclais 25 Until now: 45 rows of bricks inserted, following BAM production, 7 rows extracted to get selected bricks and reinserted In the detector: 13 rows Cside 12 rows Rside + 10 rows out for matrix code repositioning Brick filling and extraction

26 December 8 th, 2006Detector Construction status report Yves Déclais 26 Problems encountered: - high friction - CS box bent or not attached solidly - matrix code not well engraved All those points expected to be solved with next brick production Brick filling and extraction

27 December 8 th, 2006Detector Construction status report Yves Déclais 27 BMS commissioning after the run - New versions of BMM, Supervisor and operator interface - Database updated to correct bug on sign convention for detector side - Many progresses in automate programs, several bugs corrected - loss of control in vertical moves fixed - problem in inserting a 3 rd code reader found - problem of message losses after incidents understood Matrix code readers

28 December 8 th, 2006Detector Construction status report Yves Déclais 28 Mechanical interventions - Drum cage modifications on 5/6 drums: 9 th level access - Fixation of loading stations - Fixation of 4 surveying cameras (2 on each platform) - new supports for some better adapted sensors (to signal brick presence on carrousel and drums) - study of a shade for BMS lamps (with red filter towards tray) Red filter adaptation successful and adequate for avoiding spurious sensor disconnection (TT light leaks)

29 December 8 th, 2006Detector Construction status report Yves Déclais 29 Filling rate issue Until now, no effort was done to improve speed of the BMS priority was given to the security of operation. Some room for improvement – transitions -speed of individual movements Present timing for detector filling: 1 hour per row 2 shifts needed at present speed to insert 4 drums/day. Looking for factor 2 speed increase !

30 December 8 th, 2006Detector Construction status report Yves Déclais 30 Security fence Final fence ordered to BMI-Axelent firm is built and arrives next week in Gran Sasso

31 December 8 th, 2006Detector Construction status report Yves Déclais 31 Brick production restart Very crucial points for BMS: CS box fixation, final skate config. and M-code engraving Waiting presently for bricks of the future production: see Salvatore report, delivery problems Need of bricks representing the new production to perform the final BMS tuning

32 December 8 th, 2006Detector Construction status report Yves Déclais 32 Filling the detector Drum : 9 rows for a semi wall  9x26 = 234 bricks / drum Wall : 64 lines of 26 bricks per semi wall  3328 bricks / wall Target : 31 wall per Super-Module  bricks / target Filling strategy: leaving 3 rows empty at the bottom and top  3016 bricks / wall leaving 2 wall empty on both side of the target  bricks / target  ~79 % of the initial target mass Filling schedule: one SM : 27 walls, 58 lines, 116 trays  3132 trays to be filled  348 drums assuming 15 drums / week  23.2 weeks 6 months per Super Module

33 December 8 th, 2006Detector Construction status report Yves Déclais CNGS Status

34 December 8 th, 2006Detector Construction status report Yves Déclais 34 Today we have removed one shielding block of the side shielding of the CNGS reflector. The radiation was low enough to stay beside the reflector for some minutes and to have a direct look to the possible leak locations. The leak in the closed water cooling circuit of the reflector was located immediately: The insulating ceramic part of the most downstream tube connecting the outer conductor with the water drain pipe is broken (see attached picture). The reason for this rupture has to be understood. Most likely is due to additional clamps that provide vapor tightness (a condition that could be negotiated), which might overconstrain the system. These clamps could be removed quite easily, an operation that we would then like to do at all the other connecting tubes of the reflector and also the horn. We are investigating the possibility for repair of the connecting tube in-situ, i.e. without moving the reflector, but just the side shielding. In the next days we will prepare a detailled plan of the interventions during this shut-down, such that CNGS is ready for the 2007 re-start.

35 December 8 th, 2006Detector Construction status report Yves Déclais 35

36 December 8 th, 2006Detector Construction status report Yves Déclais 36

37 December 8 th, 2006Detector Construction status report Yves Déclais 37 Status of the Horn 1 uThe most important modifications concern the water cooling systems (sprayers lines and drain tubes) Helicoflex SS collar Ceramic muff with Titanium flanges TIG Welding (Al 6083) Tin/Ag O-ring SS bellow (316L) ARCLEX insulators cannot be used in wet areas Old version

38 December 8 th, 2006Detector Construction status report Yves Déclais 38 Have brazed ceramic electrical insulators on horn water lines They are strain-relieved with invar+ceramic clam-shells On horn 1, developed a 7gal/day water leak on line to spray header On horn 2, leak on suction line drew air in, reducing amount of water ejector pump could remove from horn collection tank Both ceramic sections were successfully replaced Speculate braze corrodes… planning to switch from brazed to a shrink-fit ceramic/steel connection for spare horns NUMI : Two horn water system leaks Horn 2 in February 2006, Horn 1 in August 2006

39 December 8 th, 2006Detector Construction status report Yves Déclais 39 New ceramic Water suction line NUMI Horn 2 going back in shielding after suction line replacement

40 Auxilliary drain lines Supply lines Welded bellows with ceramic isolation T2K Horn design

41 December 8 th, 2006Detector Construction status report Yves Déclais 41 Other beam aspect under investigation: Timing: Investigations with portable atomic clock. Found at CERN the reason for other 110 ns of the 600 ns of extra delay remaining. Stopped due to the beam problems. We will see how to continue in the next months at CERN without beam and,also at LNGS. Database: Claim that all the problems were fixed during the MD devoted to CNGS two weeks ago. Unfortunately we could not investigate better due to lack of beam/ We have some doubts that all the problems are solved. In the CNGS data there are from time to time missing spills and inconsistent data (e.g. protons + secondaries after the target but no muons). Padova is looking if in the data there are neutrino interactions not on-time. SPS: Stable condition with 3 CNGS cycles not reached yet, tuning interrupted by the reflector problem. Continue in the next days with dry magnet cycles or some beam to the dump

42 December 8 th, 2006Detector Construction status report Yves Déclais Perspectives for the run in 2007

43 December 8 th, 2006Detector Construction status report Yves Déclais Draft SPS schedule SPS physics run: Start: 26/5/2007 End: 7/11/ days after having excluded the MDs ( 200 days expected ) proton intensity on target canonical pot today performance pot SPSC outcome (november 06): completion of the CNGS commissioning at the beginning of the physics run no high intensity beam for the CNGS until OPERA has reached 75% of the target mass

44 December 8 th, 2006Detector Construction status report Yves Déclais 44 OPERA beam request for 2007 we should sent it before the next SPSC ( february 6-7 th )  beginning of january in the low intensity mode ( pot:extraction) we can expect: pot/day with one CNGS cycle / supercycle (18 sec) pot/day with 3 CNGS cycles / supercycle (39.6 sec) end of may we should have ~74 K bricks in the detector  6 – 9 events in the bricks per day end of october ~154 K bricks  12 – 19 events per day running over 141 days  1200 – 2000 events integrated

45 December 8 th, 2006Detector Construction status report Yves Déclais 45 Conclusions The OPERA collaboration can be proud of the real success of the construction of the detector  still one year of big efforts needed  are the needed resources available ? If we can guarantee the schedule discussed before,we can have the first physics run of OPERA next year in 2007


Download ppt "December 8 th, 2006Detector Construction status report Yves Déclais 1 OPERA Detector Preparation 1. Electronic Detectors: status of installation 2. Brick."

Similar presentations


Ads by Google